# Measuring the curvature of the universe based on the dust-scattering ring of GRB 221009A

### v1

@@ 5 Result
Table 1: $\theta$ and $\Delta t$ of scattering rings of GRB 221009A scattered by dust partly. Data is fitted by Vasilopoulos et al. [13].
rings position MJD
(arcsecond) 59862.66 59862.79 59862.87 59863.05 59863.26
$b_{\rm L,1}$ 2.466+${}_{-0.018}^{+0.017}$ 2.590+${}_{-0.026}^{+0.026}$ 2.578+${}_{-0.012}^{+0.006}$ 2.838+${}_{-0.028}^{+0.029}$ 3.116+${}_{-0.021}^{+0.022}$
$b_{\rm L,2}$ 3.342+${}_{-0.04}^{+0.029}$ 3.56+${}_{-0.05}^{+0.05}$ 3.65+${}_{-0.05}^{+0.04}$ 3.91+${}_{-0.07}^{+0.06}$ 4.18+${}_{-0.05}^{+0.05}$
$b_{\rm L,3}$ 5.700+${}_{-0.009}^{+0.009}$ 5.971+${}_{-0.017}^{+0.019}$ 6.242+${}_{-0.016}^{+0.016}$ 6.629+${}_{-0.016}^{+0.016}$ 7.115+${}_{-0.013}^{+0.014}$

### v2

@@ 5 Result
Table 1: The position of rings $\theta$ corresponds to the arrival time $T_{0}+\delta t$ of the scattering rings of GRB 221009A. We partially list three rings, $r_{1},r_{2}$ and $r_{3}$, each scattered by different dust. The data were fitted by Vasilopoulos et al. [13], and we utilized their fitted data. The unit of position is arcseconds.
rings ID MJD
59862.66 59862.79 59862.87 59863.05 59863.26
$r_{1}$ 2.466+${}_{-0.018}^{+0.017}$ 2.590+${}_{-0.026}^{+0.026}$ 2.578+${}_{-0.012}^{+0.006}$ 2.838+${}_{-0.028}^{+0.029}$ 3.116+${}_{-0.021}^{+0.022}$
$r_{2}$ 3.342+${}_{-0.04}^{+0.029}$ 3.56+${}_{-0.05}^{+0.05}$ 3.65+${}_{-0.05}^{+0.04}$ 3.91+${}_{-0.07}^{+0.06}$ 4.18+${}_{-0.05}^{+0.05}$
$r_{3}$ 5.700+${}_{-0.009}^{+0.009}$ 5.971+${}_{-0.017}^{+0.019}$ 6.242+${}_{-0.016}^{+0.016}$ 6.629+${}_{-0.016}^{+0.016}$ 7.115+${}_{-0.013}^{+0.014}$